# Nonhermitian topological zero modes at smooth domain walls: Exact solutions

### v1

@@ III.2.1 Two zero modes for $W_{L}W_{R}=-1$
For $\operatorname{sgn}({M}_{L})=\operatorname{sgn}({M}_{R})=-1$, and $\mathrm{Re}({v}_{L})\mathrm{Re}({v}_{R})<0$, the topological invariants are opposite and nonzero $W_{L}W_{R}=-1$. Moreover, in this case, one has $\operatorname{sgn}(\mathrm{Re}(\alpha_{L}^{\pm}))=-s\operatorname{sgn}(\mathrm{Re}({v}_{L}))$ and $\operatorname{sgn}(\mathrm{Re}(\alpha_{R}^{\pm}))=s\operatorname{sgn}(\mathrm{Re}({v}_{R}))$, which gives $\operatorname{sgn}(\mathrm{Re}(\alpha_{L}^{\pm}))=\operatorname{sgn}(\mathrm{Re}(\alpha_{R}^{\pm}))$ in this case. Choosing the pseudospin $s=\operatorname{sgn}{(\mathrm{Re}({v}_{L}))}=-\operatorname{sgn}{\mathrm{Re}(({v}_{R}))}$ gives $\mathrm{Re}(s{v}_{L})>0$ and $\mathrm{Re}(s{v}_{R})<0$ and therefore $\mathrm{Re}(\alpha_{L}^{\pm}),\mathrm{Re}(\alpha_{R}^{\pm})<0$: In this case, the two linearly independent solutions $\varphi_{1,2}^{s}(x)$ in Eq. 42 diverge exponentially with different decaying lengths at $x\to\pm\infty$, and their linear combination cannot satisfy the boundary conditions. Choosing instead the opposite pseudospin $s=-\operatorname{sgn}{(\mathrm{Re}({v}_{L}))}=\operatorname{sgn}{(\mathrm{Re}({v}_{R}))}$ gives $\mathrm{Re}(s{v}_{L})<0$ and $\mathrm{Re}(s{v}_{R})>0$ and therefore $\mathrm{Re}(\alpha_{L}^{\pm}),\mathrm{Re}(\alpha_{R}^{\pm})>0$: In this case, the two solutions $\varphi_{1,2}^{s}(x)$ decay exponentially to zero for $x\to\pm\infty$. Therefore, the particular solution satisfying the boundary conditions is any linear combination
@@
up to normalization constants $A_{1,2}$, with pseudospin $s=-\operatorname{sgn}{(\mathrm{Re}({v}_{L}))}=\operatorname{sgn}{(\mathrm{Re}({v}_{R}))}$ in Eq. 2, $\varphi_{1,2}^{s}(x)$ as in Eqs. 42 and 57, and ${\alpha_{L}}$, ${\alpha_{R}}$ given by Eq. 37. Since $\mathrm{Re}(\alpha_{L,R}^{-})\leq\mathrm{Re}(\alpha_{L,R}^{+})$, the asymptotic behavior at large distances is given by
@@
The asymptotic behavior of the modes $\varphi_{1}^{s}$ and $\varphi_{2}^{s}$ are determined on the left side by the exponents $\alpha_{L}^{+}=-s{v}_{L}+{q}_{L}$ (characteristic lengths $\xi_{L}^{+}$, $\lambda_{L}^{+}$) and $\alpha_{L}^{-}=-s{v}_{L}-{q}_{L}$ (characteristic lengths $\xi_{L}^{-}$, $\lambda_{L}^{-}$), respectively, and on the right side by the exponents $\alpha_{R}^{-}=s{v}_{R}-{q}_{R}$ (characteristic lengths $\xi_{R}^{-}$, $\lambda_{R}^{-}$) and $\alpha_{R}^{+}=s{v}_{R}+{q}_{R}$ (characteristic lengths $\xi_{R}^{+}$, $\lambda_{R}^{+}$), respectively.
III.2.2 One zero mode for $W_{L}=0$, $W_{R}=\pm 1$
For $\operatorname{sgn}({M}_{L})=1$, $\operatorname{sgn}({M}_{R})=-1$, and $\mathrm{Re}({v}_{R})\neq 0$, the topological invariants are $W_{L}=0$ and $W_{R}=\operatorname{sgn}(\mathrm{Re}({v}_{R}))=\pm 1$. Moreover, in this case, one has $\operatorname{sgn}(\mathrm{Re}(\alpha_{R}^{\pm}))=s\operatorname{sgn}(\mathrm{Re}({v}_{R}))$ and $\mathrm{Re}(\alpha_{L}^{+})>0>\mathrm{Re}(\alpha_{L}^{-})$. Choosing $s=-\operatorname{sgn}(\mathrm{Re}({v}_{R}))$ gives $\mathrm{Re}(\alpha_{R}^{\pm})<0$, and therefore the two linearly independent solutions $\varphi_{1,2}^{s}(x)$ in Eq. 42 respectively decay to zero and diverge for $x\to-\infty$ and both diverge for $x\to\infty$, and thus their linear combination cannot satisfy the boundary conditions. Choosing instead $s=\operatorname{sgn}(\mathrm{Re}({v}_{R}))$ gives $\mathrm{Re}(\alpha_{R}^{\pm})>0$, and therefore the solution $\varphi_{1}^{s}(x)$ now decays to zero for $x\to\pm\infty$, while $\varphi_{2}^{s}(x)$ still diverges for $x\to-\infty$. Hence, the particular solution satisfying the boundary conditions is
@@ III.2.2 One zero mode for $W_{L}=0$, $W_{R}=\pm 1$
up to a normalization constant, with pseudospin $s=\operatorname{sgn}(\mathrm{Re}({v}_{R}))$ in Eq. 2, $\varphi_{1}^{s}(x)$ as in Eqs. 42 and 57, and ${\alpha_{L}}$, ${\alpha_{R}}$ given by Eq. 37. The asymptotic behavior at large distances is given by
@@
For the limiting case $\operatorname{sgn}({M}_{L})=0$, $\operatorname{sgn}({M}_{R})=-1$, and $\mathrm{Re}({v}_{L})\mathrm{Re}({v}_{R})<0$, the solution in Eq. 16 still satisfy the boundary conditions. Indeed, by taking ${m}_{L}=0$ into the solution, one has that $\alpha_{L}^{+}>0$ if $s=-\operatorname{sgn}(\mathrm{Re}({v}_{L}))$, which is indeed the case since we choose $s=\operatorname{sgn}(\mathrm{Re}({v}_{R}))$ and $\mathrm{Re}({v}_{L})\mathrm{Re}({v}_{R})<0$.
The asymptotic behavior is determined on the left by the exponents $\alpha_{L}^{+}=-s{v}_{L}+{q}_{L}$ (characteristic lengths $\xi_{L}^{+}$, $\lambda_{L}^{+}$) and on the right by $\alpha_{R}^{-}=s{v}_{R}-{q}_{R}$ (characteristic lengths $\xi_{R}^{-}$, $\lambda_{R}^{-}$).
III.2.3 One zero mode for $W_{L}=\pm 1$, $W_{R}=0$
Analogous arguments hold in the case $\operatorname{sgn}({M}_{L})=-1$, $\operatorname{sgn}({M}_{R})=1$, and $\mathrm{Re}({v}_{L})\neq 0$, where the topological invariants are $W_{L}=\operatorname{sgn}(\mathrm{Re}({v}_{L}))=\pm 1$ and $W_{R}=0$. Moreover, in this case, one has $\operatorname{sgn}(\mathrm{Re}(\alpha_{L}^{\pm}))=-s\operatorname{sgn}(\mathrm{Re}({v}_{L}))$ and $\mathrm{Re}(\alpha_{R}^{+})>0>\mathrm{Re}(\alpha_{R}^{-})$. Choosing $s=\operatorname{sgn}(\mathrm{Re}({v}_{L}))$ gives $\mathrm{Re}(\alpha_{L}^{\pm})<0$, and therefore the two linearly independent solutions $\varphi_{1,2}^{s}(x)$ in Eq. 42 respectively diverge and decays to zero for $x\to\infty$ and both diverge for $x\to-\infty$, and thus their linear combination cannot satisfy the boundary conditions. Choosing instead $s=-\operatorname{sgn}(\mathrm{Re}({v}_{L}))$ gives $\mathrm{Re}(\alpha_{L}^{\pm})>0$, and therefore the solution $\varphi_{2}^{s}(x)$ now decays to zero for $x\to\pm\infty$ while $\varphi_{1}^{s}(x)$ diverges for $x\to\infty$. Hence, the particular solution satisfying the boundary conditions is
@@ III.2.3 One zero mode for $W_{L}=\pm 1$, $W_{R}=0$
up to a normalization constant, with corresponding pseudospin $s=-\operatorname{sgn}(\mathrm{Re}({v}_{L}))$ in Eq. 2, $\varphi_{2}^{s}(x)$ as in Eqs. 42 and 57, and ${\alpha_{L}}$, ${\alpha_{R}}$ given by Eq. 37. The asymptotic behavior at large distances is given by
@@
where $\mu_{L}^{-}=\mathrm{Re}(\alpha_{L}^{-})=-{M}_{L}$. It is clear that this case is equivalent to the previous one up to the transformation $x\to-x$, (i.e., $y\to 1-y$), which corresponds also to substituting $s\to-s$, ${\alpha_{L}}\leftrightarrow{\alpha_{R}}$, ${v}_{L}\leftrightarrow{v}_{R}$, and ${m}_{L}\leftrightarrow{m}_{R}$. This gives also ${v}_{1}\to-{v}_{1}$ and ${m}_{2}\to{m}_{2}$ in Eq. 46 which yields $a,b\to a,b$ and $1-c\to c-a-b$.
For the limiting case $\operatorname{sgn}({M}_{L})=-1$, $\operatorname{sgn}({M}_{R})=0$, and $\mathrm{Re}({v}_{L})\mathrm{Re}({v}_{R})<0$, the solution in Eq. 19 still satisfy the boundary conditions, by taking ${m}_{R}=0$ into the solution, giving ${K}_{R}^{+}>0$ if $s=\operatorname{sgn}(\mathrm{Re}({v}_{R}))$, which is indeed the case since we choose $s=-\operatorname{sgn}(\mathrm{Re}({v}_{L}))$ and $\mathrm{Re}({v}_{L})\mathrm{Re}({v}_{R})<0$.
The asymptotic behavior is determined on the left by the exponents $\alpha_{L}^{-}=-s{v}_{L}-{q}_{L}$ (characteristic lengths $\xi_{L}^{-}$, $\lambda_{L}^{-}$) and on the right by $\alpha_{R}^{+}=s{v}_{R}+{q}_{R}$ (characteristic lengths $\xi_{R}^{+}$, $\lambda_{R}^{+}$).
@@ III.2.4 No zero modes otherwise
In any other case, no nontrivial linear combination of the general solutions can satisfy the boundary conditions.
For $\operatorname{sgn}({M}_{L})\geq 0$ and $\operatorname{sgn}({M}_{R})\geq 0$, the topological invariants are zero $W_{L,R}=0$, and one has $\mathrm{Re}(\alpha_{L}^{+})\geq 0\geq\mathrm{Re}(\alpha_{L}^{-})$ and $\mathrm{Re}(\alpha_{R}^{+})\geq 0\geq\mathrm{Re}(\alpha_{R}^{-})$. The solution $\varphi_{1}^{s}(x)$ decays to zero only for $x\to-\infty$, while the solution $\varphi_{2}^{s}(y)$ decays to zero only for $x\to\infty$. Therefore, one cannot find any nontrivial linear combination that satisfies the boundary conditions.
@@
Moreover, in the case where $\operatorname{sgn}({M}_{L})=1$, $\operatorname{sgn}({M}_{R})=-1$, and $\mathrm{Re}({v}_{R})\mathrm{Re}({v}_{L})=0$, the solution Eq. 16 does not decay to zero for either $x\to\pm\infty$. Finally, in the case where $\operatorname{sgn}({M}_{L})=-1$, $\operatorname{sgn}({M}_{R})=1$, and $\mathrm{Re}({v}_{R})\mathrm{Re}({v}_{L})=0$, the solution Eq. 19 does not decay to zero for either $x\to\pm\infty$, and also in these cases, the boundary conditions cannot be satisfied.
III.3 Semi-infinite interval $0\leq x<\infty$
Zero modes on the interval $0\leq x<\infty$ satisfying the boundary conditions $\varphi(0)=\varphi(\infty)=0$ are given by linear combinations of $\varphi_{1,2}^{s}(x)$ in Eq. 7 with $\mathrm{Re}(\alpha_{R}^{\pm})>0$. As expected, there is one independent mode if $W_{R}\neq 1$. To demonstrate this statement, we need to find the particular solutions that satisfy the boundary conditions $\varphi(x\to 0)=\varphi(x\to\infty)=0$, i.e., $\varphi(y\to 1/2)=\varphi(y\to 1)=0$.
Assuming $|F_{1,2}(y)|$ convergent or logarithmically divergent at $y=1/2,1$, we only need to care now about the factor $(1-y)^{w\alpha_{R}^{-}}$, which converge to zero at $y=1$ only if $\mathrm{Re}({\alpha_{R}^{-}})>0$. This gives only two possible outcomes, discussed below and summarized in Table 2. Zero modes on the interval $-\infty<x\leq 0$ can be obtained similarly. Note that the confinement to semi-infinite intervals introduces a sharp domain wall at $x=0$, being effectively equivalent to an infinite potential wall and which is superimposed to a smooth domain wall with a finite width when $w>0$.

### v2

@@ III.2.1 Two zero modes for $W_{L}W_{R}=-1$
For $\operatorname{sgn}({M}_{L})=\operatorname{sgn}({M}_{R})=-1$, and $\mathrm{Re}({v}_{L})\mathrm{Re}({v}_{R})<0$, the topological invariants are opposite and nonzero $W_{L}W_{R}=-1$. Moreover, in this case, one has $\operatorname{sgn}(\mathrm{Re}(\alpha_{L}^{\pm}))=-s\operatorname{sgn}(\mathrm{Re}({v}_{L}))$ and $\operatorname{sgn}(\mathrm{Re}(\alpha_{R}^{\pm}))=s\operatorname{sgn}(\mathrm{Re}({v}_{R}))$, which gives $\operatorname{sgn}(\mathrm{Re}(\alpha_{L}^{\pm}))=\operatorname{sgn}(\mathrm{Re}(\alpha_{R}^{\pm}))$ in this case. Choosing the pseudospin $s=\operatorname{sgn}{(\mathrm{Re}({v}_{L}))}=-\operatorname{sgn}{\mathrm{Re}(({v}_{R}))}$ gives $\mathrm{Re}(s{v}_{L})>0$ and $\mathrm{Re}(s{v}_{R})<0$ and therefore $\mathrm{Re}(\alpha_{L}^{\pm})$ and $\mathrm{Re}(\alpha_{R}^{\pm})<0$: In this case, the two linearly independent solutions $\varphi_{1,2}^{s}(x)$ in Eq.˜42 diverge exponentially with different decaying lengths at $x\to\pm\infty$, and their linear combination cannot satisfy the boundary conditions. Choosing instead the opposite pseudospin $s=-\operatorname{sgn}{(\mathrm{Re}({v}_{L}))}=\operatorname{sgn}{(\mathrm{Re}({v}_{R}))}$ gives $\mathrm{Re}(s{v}_{L})<0$ and $\mathrm{Re}(s{v}_{R})>0$ and therefore $\mathrm{Re}(\alpha_{L}^{\pm})$ and $\mathrm{Re}(\alpha_{R}^{\pm})>0$: In this case, the two solutions $\varphi_{1,2}^{s}(x)$ decay exponentially for $x\to\pm\infty$. Therefore, the eigenmodes satisfying the boundary conditions are
@@
up to normalization constants $A_{1,2}$, with pseudospin $s=-\operatorname{sgn}{(\mathrm{Re}({v}_{L}))}=\operatorname{sgn}{(\mathrm{Re}({v}_{R}))}$ and $\varphi_{1,2}^{s}(x)$ as in Eqs.˜42 and 57, and $\alpha_{L}$}, $\alpha_{R}$} given by Eq.˜37. Since $\mathrm{Re}(\alpha_{L,R}^{-})\leq\mathrm{Re}(\alpha_{L,R}^{+})$, the asymptotic behavior is
@@
The asymptotic behavior of the two modes $\varphi_{1}^{s}$ and $\varphi_{2}^{s}$ on the left of the smooth domain wall is determined by the exponents $\alpha_{L}^{+}=-s{v}_{L}+{q}_{L}$ (decay length $\xi_{L}^{+}$ and oscillation wavelength $\lambda_{L}^{+}$) and $\alpha_{L}^{-}=-s{v}_{L}-{q}_{L}$ (decay length $\xi_{L}^{-}$ and oscillation wavelength $\lambda_{L}^{-}$), respectively, and on the right by the exponents $\alpha_{R}^{-}=s{v}_{R}-{q}_{R}$ (decay length $\xi_{R}^{-}$ and oscillation wavelength $\lambda_{R}^{-}$) and $\alpha_{R}^{+}=s{v}_{R}+{q}_{R}$ (decay length $\xi_{R}^{+}$ and oscillation wavelength $\lambda_{R}^{+}$), respectively.
III.2.2 One zero mode for $W_{L}=0$ and $|W_{R}|=1$
For $\operatorname{sgn}({M}_{L})=1$, $\operatorname{sgn}({M}_{R})=-1$, and $\mathrm{Re}({v}_{R})\neq 0$, the topological invariants are $W_{L}=0$ and $W_{R}=\operatorname{sgn}(\mathrm{Re}({v}_{R}))=\pm 1$. Moreover, in this case, one has $\operatorname{sgn}(\mathrm{Re}(\alpha_{R}^{\pm}))=s\operatorname{sgn}(\mathrm{Re}({v}_{R}))$ and $\mathrm{Re}(\alpha_{L}^{+})>0>\mathrm{Re}(\alpha_{L}^{-})$. Choosing $s=-\operatorname{sgn}(\mathrm{Re}({v}_{R}))$ gives $\mathrm{Re}(\alpha_{R}^{\pm})<0$, and therefore the two linearly independent solutions $\varphi_{1,2}^{s}(x)$ in Eq.˜42 respectively decay and diverge for $x\to$-$\infty$ and both diverge for $x\to\infty$, and thus their linear combination cannot satisfy the boundary conditions. Choosing instead $s=\operatorname{sgn}(\mathrm{Re}({v}_{R}))$ gives $\mathrm{Re}(\alpha_{R}^{\pm})>0$, and consequently the solution $\varphi_{1}^{s}(x)$ now decays for $x\to\pm\infty$, while $\varphi_{2}^{s}(x)$ still diverges for $x\to$-$\infty$. Hence, the eigenmode satisfying the boundary conditions is
@@
up to a normalization constant, with pseudospin $s=\operatorname{sgn}(\mathrm{Re}({v}_{R}))$ and $\varphi_{1}^{s}(x)$ as in Eqs.˜42 and 57, and $\alpha_{L}$}, $\alpha_{R}$} given by Eq.˜37. The asymptotic behavior is
@@
For the limiting case $\operatorname{sgn}({M}_{L})=0$, $\operatorname{sgn}({M}_{R})=-1$, and $\mathrm{Re}({v}_{L})\mathrm{Re}({v}_{R})<0$, the solution in Eq.˜16 still satisfy the boundary conditions. Indeed, by taking ${m}_{L}=0$ into the solution, one has that $\alpha_{L}^{+}>0$ if $s=-\operatorname{sgn}(\mathrm{Re}({v}_{L}))$, which is indeed the case since we choose $s=\operatorname{sgn}(\mathrm{Re}({v}_{R}))$ and $\mathrm{Re}({v}_{L})\mathrm{Re}({v}_{R})<0$.
The asymptotic behavior on the left is determined by the exponents $\alpha_{L}^{+}=-s{v}_{L}+{q}_{L}$ (decay length $\xi_{L}^{+}$ and oscillation wavelength $\lambda_{L}^{+}$) and on the right by $\alpha_{R}^{-}=s{v}_{R}-{q}_{R}$ (decay length $\xi_{R}^{-}$ and oscillation wavelength $\lambda_{R}^{-}$).
III.2.3 One zero mode for $|W_{L}|=1$, $W_{R}=0$
Analogous arguments hold if $\operatorname{sgn}({M}_{L})=-1$, $\operatorname{sgn}({M}_{R})=1$, and $\mathrm{Re}({v}_{L})\neq 0$, where the topological invariants are $W_{L}=\operatorname{sgn}(\mathrm{Re}({v}_{L}))=\pm 1$ and $W_{R}=0$. Moreover, in this case, one has $\operatorname{sgn}(\mathrm{Re}(\alpha_{L}^{\pm}))=-s\operatorname{sgn}(\mathrm{Re}({v}_{L}))$ and $\mathrm{Re}(\alpha_{R}^{+})>0>\mathrm{Re}(\alpha_{R}^{-})$. Choosing $s=\operatorname{sgn}(\mathrm{Re}({v}_{L}))$ yields $\mathrm{Re}(\alpha_{L}^{\pm})<0$, and therefore the two linearly independent solutions $\varphi_{1,2}^{s}(x)$ in Eq.˜42 respectively diverge and decays for $x\to\infty$ and both diverge for $x\to$-$\infty$, and thus their linear combination cannot satisfy the boundary conditions. Choosing instead $s=-\operatorname{sgn}(\mathrm{Re}({v}_{L}))$ gives $\mathrm{Re}(\alpha_{L}^{\pm})>0$, and therefore the solution $\varphi_{2}^{s}(x)$ now decays for $x\to\pm\infty$ while $\varphi_{1}^{s}(x)$ diverges for $x\to\infty$. Hence, the eigenmode satisfying the boundary conditions is
@@
up to a normalization constant, with corresponding pseudospin $s=-\operatorname{sgn}(\mathrm{Re}({v}_{L}))$ and $\varphi_{2}^{s}(x)$ as in Eqs.˜42 and 57, and $\alpha_{L}$}, $\alpha_{R}$} given by Eq.˜37. The asymptotic behavior is
@@
where $\mu_{L}^{-}=\mathrm{Re}(\alpha_{L}^{-})=-{M}_{L}$. It is clear that this case is equivalent to the previous one up to the transformation $x\to-x$, (i.e., $y\to 1-y$), which corresponds also to substituting $s\to-s$, ${\alpha_{L}}\leftrightarrow{\alpha_{R}}$, ${v}_{L}\leftrightarrow{v}_{R}$, and ${m}_{L}\leftrightarrow{m}_{R}$. This gives also ${v}_{1}\to-{v}_{1}$ and ${m}_{2}\to{m}_{2}$ in Eq.˜46 which yields $a,b\to a,b$ and $1-c\to c-a-b$.
For the limiting case $\operatorname{sgn}({M}_{L})=-1$, $\operatorname{sgn}({M}_{R})=0$, and $\mathrm{Re}({v}_{L})\mathrm{Re}({v}_{R})<0$, the solution in Eq.˜19 still satisfy the boundary conditions, by taking ${m}_{R}=0$ into the solution, giving ${K}_{R}^{+}>0$ if $s=\operatorname{sgn}(\mathrm{Re}({v}_{R}))$, which is indeed the case since we choose $s=-\operatorname{sgn}(\mathrm{Re}({v}_{L}))$ and $\mathrm{Re}({v}_{L})\mathrm{Re}({v}_{R})<0$.
The asymptotic behavior on the left is determined by the exponents $\alpha_{L}^{-}=-s{v}_{L}-{q}_{L}$ (decay length $\xi_{L}^{-}$ and oscillation wavelength $\lambda_{L}^{-}$) and on the right by $\alpha_{R}^{+}=s{v}_{R}+{q}_{R}$ (decay length $\xi_{R}^{+}$ and oscillation wavelength $\lambda_{R}^{+}$).
@@ III.2.4 No zero modes otherwise
In any other case, no linear combination of the general solutions satisfies the boundary conditions with the exception of the trivial solution $\varphi(x)=0$.
For $\operatorname{sgn}({M}_{L})\geq 0$ and $\operatorname{sgn}({M}_{R})\geq 0$, the topological invariants are zero $W_{L,R}=0$, and one has $\mathrm{Re}(\alpha_{L}^{+})\geq 0\geq\mathrm{Re}(\alpha_{L}^{-})$ and $\mathrm{Re}(\alpha_{R}^{+})\geq 0\geq\mathrm{Re}(\alpha_{R}^{-})$. The solution $\varphi_{1}^{s}(x)$ decays only for $x\to$-$\infty$, while the solution $\varphi_{2}^{s}(y)$ decays only for $x\to\infty$. Therefore, one cannot find any nontrivial linear combination that satisfies the boundary conditions.
@@
Moreover, in the case where $\operatorname{sgn}({M}_{L})=1$, $\operatorname{sgn}({M}_{R})=-1$, and $\mathrm{Re}({v}_{R})\mathrm{Re}({v}_{L})=0$, the solution Eq.˜16 does not decay for either $x\to\pm\infty$. Finally, in the case where $\operatorname{sgn}({M}_{L})=-1$, $\operatorname{sgn}({M}_{R})=1$, and $\mathrm{Re}({v}_{R})\mathrm{Re}({v}_{L})=0$, the solution Eq.˜19 does not decay for either $x\to\pm\infty$, and also in these cases, the boundary conditions cannot be satisfied.
III.3 Semi-infinite interval $0,\infty)$
We now consider zero-energy modes in a semi-infinite system. These are defined as the eigenmodes with zero energy that satisfy the boundary conditions $\varphi(0)=\varphi(\infty)=0$ on the interval $0,\infty)$. In order to satisfy the boundary conditions, the modes need to decay asymptotically for $x\to\infty$, which mandates that $\mathrm{Re}(\alpha_{R}^{\pm})>0$. As expected, there is one independent mode if $W_{R}\neq 1$. To demonstrate this statement, we need to find the particular solutions that satisfy the boundary conditions $\varphi(x\to 0)=\varphi(x\to\infty)=0$, i.e., $\varphi(y\to 1/2)=\varphi(y\to 1)=0$.
Assuming $|F_{1,2}(y)|$ convergent or logarithmically divergent at $y=1/2,1$, we only need to care now about the factor $(1-y)^{w\alpha_{R}^{-}}$, which converge to zero at $y=1$ only if $\mathrm{Re}({\alpha_{R}^{-}})>0$. This gives only two possible outcomes, discussed below and summarized in Table˜2. Zero modes on the interval $(-\infty,0]$ are obtained similarly. Note that the confinement to semi-infinite intervals introduces a sharp domain wall at $x=0$, being effectively equivalent to an infinite potential wall superimposed to a smooth domain wall with a finite width when $w>0$.